# Hidden heavy flavor tetraquarks in the Born-Oppenheimer approximation

### v1

@@ II.3 Wave function
Now we construct the basis vector wave functions of the tetraquark state in spin space. And the total spin of all possible $S$-wave tetraquark states be 0, 1, and 2. For $J^{P(C)}=0^{+(+)}$, there are two spin-zero states,
$|0^{+(+)}\rangle$ $=|0_{Qq},0_{\bar{Q}\bar{q}};J=0\rangle,$ (30)
$|0^{+(+)}\rangle$ $=|1_{Qq},1_{\bar{Q}\bar{q}};J=0\rangle.$
The (+/)(+/-)( + / - ) represents charge conjugation when the tetraquark state is in flavor symmetry, and has the same meaning as below unless otherwise specified.
For $J=1$ with positive parity, there are three states,
@@
Under charge conjugation, $|A\rangle$ and $|B\rangle$ are exchanged, while $|C\rangle$ has odd parity. Therefore, the $1^{+}$ composite state consists of one $C$-even state and two $C$-odd states,
@@
$|1^{+(-)}\rangle$ $=\frac{1}{\sqrt{2}}(|A\rangle-|B\rangle),$
$|1^{+(-)^{\prime}}\rangle$ $=|C\rangle.$
We can analyze the specific spin values of these states for both $Q\bar{Q}$ and $q\bar{q}$ quark pairs. Since $J=1$, it is impossible to have a state with zero spin; among the remaining states, only the $C=+$ state has spin 1 for both quarks. Therefore, the state $|1^{+(+)}\rangle$ in Eq.(32) has a definite $Q\bar{Q}$ spin value of $S_{Q\bar{Q}}=1$. For $J^{PC}=2^{+(+)}$,
@@
The $2^{+(+)}$ state also has $S_{Q\bar{Q}}=1$. Therefore, in order to solve the mass spectra and wave functions of the tetraquark states, we only need to substitute the basis vectors into Eq.(2) and diagonalize them.
@@ III.1 Model parameters
In numerical calculations, we take the values of $\alpha_{S}$ and $\sigma$, based on the charmonium scale obtained from lattice calculations Kawanai and Sasaki (2012),
$\alpha_{S}(2M_{c})=0.30,\quad\sigma=0.15\,\text{GeV}^{2}.$ (34)
At the $B_{c}$ meson and bottom-charm quark mass scale, we use the same string tension $\sigma$ and $\alpha_{S}$ running via the two-loop beta function, yielding
$\alpha_{S}(M_{c}+M_{b})=0.24,\quad\alpha_{S}(2M_{b})=0.21.$ (35)
Table 1: Constituent quark masses derived from the $S$-wave mesons (first row) and the baryons (second row).
$q$ $s$ $c$ $b$
Mass1 (MeV) 305 490 1670 5005
Mass2 (MeV) 362 546 1721 5044
Table 2: Spin-spin couplings for quark-antiquark pairs in color singlet from the hyperfine splittings of $S$-wave mesons (first row). The values in the second row show the approximate scaling of the couplings with inverse masses (from the meson spectra).
$q\bar{q}$ $s\bar{q}$ $s\bar{s}$ $c\bar{q}$ $c\bar{s}$ $c\bar{c}$
$(\kappa_{ij})_{\bm{0}}$ (MeV) 315 195 121 70 72 59
$(\kappa_{ij})_{\bm{0}}m_{i}m_{j}$ (GeV3) 0.029 0.029 0.036 0.059 0.16
Table 3: Spin-spin couplings for quark-quark pairs in color $\bar{\bm{3}}$ state from $S-$wave baryons. One gluon exchange implies $(\kappa_{ij})_{\bar{\bm{3}}}=1/2(\kappa_{ij})_{\textbf{0}}$. The values in the second row show the approximate scaling of the couplings with inverse masses (from the baryon spectra).
$qq$ $sq$ $cq$ $cs$
$(\kappa_{ij})_{\bar{\bm{3}}}$ (MeV) 103 64 22 25
$(\kappa_{ij})_{\bar{\bm{3}}}m_{i}m_{j}$ (GeV3) 0.014 0.013 0.014 0.024
The numerical solution of the Schrödinger equation will yield the energy eigenvalues $E$ of the tetraquark wave function. However, there still exists a constant $V_{0}$ in Eq. (20) that influences the exact value of the eigenvalue. For tetraquark states, we can approximate the boundary conditions to derive the relationship between $V_{0}$ and $\langle H\rangle_{\text{min}}$, but for hidden heavy flavor tetraquarks states, the boundary conditions are difficult to resolve. We trace back to the original meaning of $V_{0}$, which originates from the Cornell potential as a fitting constant in both experimental and theoretical models. In our analysis, we utilize the $X(3872)$ as an input parameter for the $c\bar{c}q\bar{q}$ system. Through this approach, we determine the optimal solution for $V_{0}=-0.315$ GeV. This value is consistent with the Cornell potential model. We can write the complete expression for the $c\bar{c}q\bar{q}$ mass as
$M(T)=2(M_{c}+M_{q})+E+H_{cm},$ (36)
and plot the relationship between $M(T)$ and $R_{0}$, see Fig.1.
Figure 1: For $J^{PC}=1^{++}$, the ground state energy of the $c\bar{c}q\bar{q}$ system decreases as the system size $R_{0}$ increases.
Besides, we calculated the mass spectra of the tetraquark state at different $R_{0}$, see Table 4. Among them, when $R_{0}=10$ GeV-1, the mass of the tetraquark state $c\bar{c}q\bar{q}$ with quantum number $1^{++}$ is 3.872 GeV.
Table 4: Mass spectra of tetraquark $c\bar{c}q\bar{q}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
@@
$R_{0}=6$ 3.792, 3.898 3.785, 3.913 3.903 3.982
$R_{0}=8$ 3.770, 3.876 3.763, 3.891 3.881 3.960
$R_{0}=10$ 3.761, 3.867 3.754, 3.882 3.872 3.951
$R_{0}=12$ 3.757, 3.863 3.750, 3.878 3.868 3.947
Using this method, we can estimate the mass spectra of other tetraquark states.
Figure 2: The mass spectra of the $S$-wave tetraquark states with distinct quantum numbers: (a) $c\bar{c}q\bar{q}$, (b) $c\bar{c}q\bar{w}$, and (c) $c\bar{c}s\bar{s}$. The black dashed lines and red dash-dotted lines represent the corresponding meson-meson thresholds.
Figure 3: The mass spectra of the $S$-wave tetraquark states with distinct quantum numbers: (a) $b\bar{b}q\bar{q}$, (b) $b\bar{b}q\bar{s}$, and (c) $b\bar{b}s\bar{s}$. The black dashed lines and red dash-dotted lines represent all possible meson-meson thresholds.
III.2 Hidden-charm tetraquark states with $c\bar{c}q\bar{q}$, $c\bar{c}q\bar{s}$, and $c\bar{c}s\bar{s}$ configurations
III.2.1 The $c\bar{c}q\bar{q}$ configuration
As long as $Q(q)$ and $\bar{Q}(\bar{q})$ are of the same flavor, the tetraquark state has a well-defined $C$-parity. The tetraquark state can strongly decay into a pair of mesons, and in these processes, the $C$-parity is conserved.
(1) The mass spectrum of $c\bar{c}q\bar{q}$ is shown in FIG. 2(a), with an initial state having $J^{P}=0^{+}$ ($J^{PC}=0^{++}$), which is electrically neutral. Their masses are 3761.1 MeV and 3867.2 MeV, respectively. The lowest state, at 3761.1 MeV, is above the $D\bar{D}$ threshold and can decay via S-wave into $D\bar{D}$ and $\eta_{c}\pi$. The state at 3867.2 MeV is also above the $\eta_{c}\pi$ and $D\bar{D}$ thresholds but slightly below the $J/\psi\rho$ threshold. Considering angular momentum and $C$-parity conservation, this state allows strong decays into $\eta_{c}\pi$, $D\bar{D}$, or $J/\psi\rho$ mesons. This state could be a strong contender for the $X(3915)$ candidate. In this paper, we focus only on $S$-wave decays and neglect higher-order decays, such as $P$-wave and $D$-wave, as these typically contribute less.
(2) There are three $S$-wave tetraquark states with $J^{P}=1^{+}$. The electrically neutral tetraquark states can be classified into two categories based on their $C$-parity. Two $S$-wave tetraquark states have $J^{PC}=1^{+-}$. The lowest state, with a mass of 3754.1 MeV, is slightly below the $\eta_{c}\rho$ threshold. It can easily decay into $\eta_{c}$ and $\rho$, as it has a significant component and a small mass difference. Similarly, the state with a mass of 3882.3 MeV is close to the $D\bar{D}^{*}$ threshold, suggesting that it can easily decay into $D\bar{D}^{*}$. Additionally, it can also decay into $\eta_{c}\rho$ or $J/\psi\pi$. It’s encouraging that these decay channels are considered as potential decay routes for $X(3900)$, which strongly supports the candidacy of this state as $X(3900)$. Another $S$-wave tetraquark state with positive $C$-parity is the one with $J^{PC}=1^{++}$. Its mass is 3872 MeV, which is the result we obtained by fitting the experimental value X(3872). In our model, the most likely decay mode for this particle is $J/\psi\rho$, a result consistent with the findings in Braaten and Kusunoki (2005).
(3) There is an $S$-wave tetraquark state with $J^{P}=2^{+}$ ($J^{PC}=2^{++}$) in the electrically neutral state. Its mass is 3951.1 MeV, which is above the $J/\psi\rho$ threshold, making it possible for the state to decay into $J/\psi\rho$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in Table 4, and the decay analysis method remains consistent with the aforementioned approach.
III.2.2 The $c\bar{c}q\bar{s}$ configuration
The mass spectrum of $c\bar{c}q\bar{s}$ is shown in FIG. 2(b). For comparison, the corresponding meson-meson thresholds are also presented in the figure. For the tetraquark state of $c\bar{c}q\bar{s}$, the $C$-parity is undetermined, so the quantum numbers of the $S$-wave tetraquark state $c\bar{c}q\bar{s}$ are $J^{P}=0^{+}$, $1^{+}$, and $2^{+}$.
(1) There are two $J^{P}=0^{+}$ states, with masses of 3901.1 MeV and 3983.3 MeV, respectively. The lighter state, at 3901.1 MeV, is above the $\eta_{c}K$ and $D\bar{D}_{s}$ thresholds, and can easily decay into $\eta_{c}K$ and $D\bar{D}_{s}$. The state at 3983.3 MeV not only decays naturally into the mesons $\eta_{c}K$ and $D\bar{D}_{s}$, but is also very close to the $J/\psi K^{*}$ threshold, allowing for the possibility of decaying into the meson pair $J/\psi K^{*}$.
(2) There are three $S$-wave tetraquark states with $J^{P}=1^{+}$. Their masses are 3913.1, 3970, and 3983.2 MeV, respectively. The lightest state, with a mass of 3913.1 MeV, is above the thresholds for $J/\psi K$ and $\eta_{c}K^{*}$, and its decay channels are likely to include $J/\psi K$ or $\eta_{c}K^{*}$. Similarly, the state at 3970 MeV also has decay channels that may include $J/\psi K$ or $\eta_{c}K^{*}$. The state at 3983.2 MeV, which lies just below the $J/\psi K^{*}$ threshold and slightly above the $D\bar{D}_{s}^{*}$ threshold, is most likely to decay into $J/\psi K$, $\eta_{c}K^{*}$, $J/\psi K^{*}$, or $D\bar{D}_{s}^{*}$. It is worth mentioning that the states at 3970 and 3983.2 MeV could potentially be candidates for $Z_{cs}(3985)^{-}$ and $Z_{cs}(4000)^{+}$.
(3) For $J^{P}=2^{+}$, there is an $S$-wave tetraquark state with a mass of 3988.9 MeV. Its mass is slightly below the threshold for $D^{*}\bar{D}_{s}^{*}$, allowing it to naturally decay into the meson pairs $J/\psi K$, $\eta_{c}K^{*}$, $J/\psi K^{*}$, or $D\bar{D}_{s}^{*}$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in TABLE 5, and the decay analysis method remains consistent with the aforementioned approach.
Table 5: Mass spectra of tetraquark $c\bar{c}q\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)
@@ III.2.2 The $c\bar{c}q\bar{s}$ configuration
$R_{0}=6$ 3.929, 4.011 3.942, 3.998, 4.011 4.017
$R_{0}=8$ 3.909, 3.991 3.922, 3.978, 3.991 3.997
$R_{0}=10$ 3.901, 3.983 3.913, 3.970, 3.983 3.989
$R_{0}=12$ 3.898, 3.980 3.910, 3.967, 3.979 3.986
III.2.3 The $c\bar{c}s\bar{s}$ configuration
The tetraquark configuration $c\bar{c}s\bar{s}$ has a well-defined $C$-parity. The possible quantum numbers are $J^{PC}=0^{++},1^{+-},2^{++}$. The mass spectrum of the tetraquark states with the $c\bar{s}c\bar{s}$ configuration is shown in FIG. 2(c), along with the corresponding meson-meson thresholds.
(1) There are two $S$-wave tetraquark states with $J^{PC}=0^{++}$, having masses of 4032.2 MeV and 4122.3 MeV, respectively. The lowest mass state, at 4032.2 MeV, is above the $D_{s}\bar{D}_{s}$ meson-meson threshold, so its possible decay mode is into $D_{s}\bar{D}_{s}$. The state at 4122.3 MeV, which lies just below the $D_{s}^{*}\bar{D}_{s}^{*}$ threshold, can naturally decay into $D_{s}\bar{D}_{s}$ and $J/\psi\phi$.
(2) There are two $S$-wave tetraquark states with $J^{PC}=1^{+-}$, having masses of 4069.1 MeV and 4102.8 MeV, respectively. The lighter state, at 4069.1 MeV, is well above the $\eta_{c}\phi$ threshold, allowing it to naturally decay into $\eta_{c}\phi$. The state at 4102.8 MeV, in addition to the decay mode into $\eta_{c}\phi$, also has a possible decay mode into $D_{s}\bar{D}_{s}^{*}$. Furthermore, there is an $S$-wave tetraquark state with $J^{PC}=1^{++}$ and a mass of 4088.2 MeV. For this state, both $\eta_{c}\phi$ and $D_{s}\bar{D}_{s}^{*}$ decay modes are allowed. Although several promising exotic $c\bar{c}s\bar{s}$ tetraquark candidates had been identified prior to this, our experimental theoretical results more strongly support $X(4140)$ as a $c\bar{c}s\bar{s}$ tetraquark state.
(3) For $J^{PC}=2^{++}$, we observe an $S$-wave tetraquark state with a mass of 4174.3 MeV. Considering $J^{PC}$ conservation, this state can only decay into the meson pair $J/\psi\phi$. Additionally, we have calculated the masses of tetraquarks with varying sizes, as depicted in Table 6, and the decay analysis method remains consistent with the aforementioned approach.
Table 6: Mass spectra of tetraquark $c\bar{c}s\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (GeV)

### v2

@@ II.3 Wave function
In order to study the energy level splitting caused by color-spin hyperfine interaction, we can construct the basis vector wave functions of the tetraquark state. In color space, the tetraquark state is divided into meson-meson configuration $|(Q\bar{Q})(q\bar{q})\rangle$ (or $|(Q\bar{q})(q\bar{Q})\rangle$) and diquark-antidiquark configuration $|(Qq)(\bar{Q}\bar{q})\rangle$. And these two configurations can be connected by Fierz transformation. In this calculation, we focus only on the diquark-antidiquark configuration. Obviously, the total spin of all possible $S$-wave tetraquark states be 0, 1, and 2. For the scalar tetraquark states with quantum number $J^{P}=0^{+}$, there are two basis vectors
$|0^{+(+)}\rangle_{1}=|0_{Qq},0_{\bar{Q}\bar{q}};J=0\rangle,$ (24)
$|0^{+(+)}\rangle_{2}=|1_{Qq},1_{\bar{Q}\bar{q}};J=0\rangle.$
The symbol (+/)(+/-)( + / - ) represents charge conjugation when a tetraquark state is in flavor symmetry. And the positive and negative signs in quantum numbers have the same meaning below, unless otherwise stated.
For the axial vector tetraquark states with quantum number $J^{P}=1^{+}$, there are three basis vectors
@@
In charge conjugation, considering the symmetric and antisymmetric combinations of states $|A\rangle$ and $|B\rangle$, we obtain a $C$-odd and a $C$-even state. And the charge conjugation of $|C\rangle$ is odd. Therefore, for the hidden heavy flavor tetraquark states with $J^{P}=1^{+}$, the basis vectors with definite charge conjugation can be rewritten as
@@
$|1^{+(-)}\rangle_{1}=\frac{1}{\sqrt{2}}(|A\rangle-|B\rangle),$
$|1^{+(-)}\rangle_{2}=|C\rangle.$
For the $J^{P}=2^{+}$ state, the basis vector is
@@
Now we have constructed the basis vector wave functions of the hidden heavy flavor tetraquark system based on the diquark-antidiquark configuration. Substituting these basis vectors into Eq.(2), we can obtain the corresponding Hamiltonian matrices. Then, by diagonalizing the Hamiltonian matrices, we can obtain the mass spectra and corresponding wave functions of the tetraquark states. Finally, considering the characteristics of the meson-meson configuration, we analyze possible decay channels of these calculated tetraquark states.
@@ III.1 Model parameters
In the specific numerical calculation, we first need to determine the model parameters. By fitting the experimental observations and lattice calculations of heavy flavor conventional hadrons, we extract the effective masses of the constituent quarks: $m_{n}=308$ MeV, $m_{s}=484$ MeV, $m_{c}=1667$ MeV, and $m_{b}=5005$ MeV. Here $n$ represents $u$ and $d$ quarks. Without loss of generality, we take $\alpha_{s}$ and $\sigma$ based on the charmonium, bottomonium, and bottom-charm mesons: $\alpha_{s}(2m_{c})=0.30$, $\alpha_{s}(m_{c}+m_{b})=0.24$, $\alpha_{s}(2m_{b})=0.21$, and $\sigma=0.15$ GeV2. In addition, other model parameters required for our calculation are listed in the Tables 1-2.
Table 1: Parameter $\kappa$ (in MeV) for quark-antiquark systems. And these parameters are derived from fitting the $S$-wave mesons.
$n\bar{n}$ $s\bar{n}$ $s\bar{s}$ $c\bar{n}$ $c\bar{s}$ $c\bar{c}$ $b\bar{n}$ $b\bar{s}$ $b\bar{c}$ $b\bar{b}$
$(\kappa_{ij})_{\bm{0}}$ 315 195 121 70 72 59 23 24 20 30
Table 2: Parameter $\kappa$ (in MeV) for quark-quark (or antiquark-antiquark) systems. And these parameters are derived from fitting the $S$-wave baryons.
$nn$ $sn$ $cn$ $cs$ $bn$ $bs$
$(\kappa_{ij})_{\bar{\bm{3}}}$ 103 64 22 25 6.6 7.5
So far, since the $s\bar{s}$ meson with quantum number $J^{P}=0^{-}$ and the $b\bar{c}$ meson with quantum number $J^{P}=1^{-}$ have not been observed experimentally, the parameters $\kappa_{s\bar{s}}$ and $\kappa_{b\bar{c}}$ can be extracted by using a similar method as in Ref. Maiani et al. (2005). It is worth emphasizing that these parameters have a certain symmetry, namely $\kappa_{ij}=\kappa_{ji}$.
Figure 1: For the tetraquark state $c\bar{c}n\bar{n}$ with quantum number $J^{PC}=1^{++}$, the mass spectrum of the ground state decreases with increasing parameter $R_{0}$.
To solve the mass spectra of the hidden heavy flavor tetraquark states, we choose $X(3872)$ as input to determine $R_{0}$. Assuming that $X(3872)$ is a tetraquark state $c\bar{c}n\bar{n}$ with quantum number $J^{PC}=1^{++}$, we can substitute the basis vector $|1^{++}\rangle$ into the Hamiltonian Eq. (2) to obtain the corresponding mass spectrum
$M(T)=2(m_{c}+m_{q})+E+H_{ss},$ (28)
where $T$ represents the tetraquark state $c\bar{c}n\bar{n}$. Now we plot the mass $M(T)$ as $R_{0}$ changes, see Figure 1. It can be found that the mass $M(T)$ decreases as $R_{0}$ increases. It is noteworthy that when $R_{0}\geq$ 10 $\text{GeV}^{-1}$, the mass spectrum of the tetraquark state $c\bar{c}n\bar{n}$ tends to be constant. This indicates that the confinement between $c\bar{c}$ disappears and the tetraquark state decays into two open-charm mesons. Besides, we calculated the mass spectra of the tetraquark state $c\bar{c}n\bar{n}$ at different $R_{0}$, see Table 3. Among them, when $R_{0}=10\,\text{GeV}^{-1}$, the mass spectrum of the tetraquark state $c\bar{c}n\bar{n}$ with quantum number $1^{++}$ is 3872 MeV.
Table 3: Mass spectra of tetraquark $c\bar{c}n\bar{n}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)
$M(0^{+(+)})$ $M(1^{+(-)})$ $M(1^{+(+)})$ $M(2^{+(+)})$
$R_{0}=6$ 3792, 3898 3785, 3913 3903 3982
$R_{0}=8$ 3770, 3876 3763, 3891 3881 3960
$R_{0}=10$ 3761, 3867 3754, 3882 3872 3951
$R_{0}=12$ 3757, 3863 3750, 3878 3868 3947
Figure 2: The mass spectra of the $S$-wave hidden-charm tetraquark states: (a) $c\bar{c}n\bar{n}$, (b) $c\bar{c}n\bar{s}$, and (c) $c\bar{c}s\bar{s}$. The black dashed lines and red dash-dotted lines represent the corresponding meson-meson thresholds.
Figure 3: The mass spectra of the $S$-wave hidden-bottom tetraquark states: (a) $b\bar{b}n\bar{n}$, (b) $b\bar{b}n\bar{s}$, and (c) $b\bar{b}s\bar{s}$. The black dashed lines and red dash-dotted lines represent all possible meson-meson thresholds.
III.2 Hidden-charm tetraquark states with $c\bar{c}n\bar{n}$, $c\bar{c}n\bar{s}$, and $c\bar{c}s\bar{s}$ configurations
In this section, we systematically investigate the mass spectra of the $S$-wave hidden-charm tetraquark states by properly considering the BO approximation and the color-spin hyperfine interaction. And the quantum number assignments of these tetraquark states are $J^{P(C)}=0^{+(+)}$, $1^{+(\pm)}$, and $2^{+(+)}$. The model parameters and specific calculation scheme have been explained in the previous section. We take $R_{0}=10$ GeV-1, and the mass spectra of the tetraquark states $c\bar{c}n\bar{n}$, $c\bar{c}n\bar{s}$, and $c\bar{c}s\bar{s}$ are plotted in Figure 2, respectively. For ease of comparison, the corresponding meson-meson thresholds are also plotted in these subfigures. Besides, we have also calculated the mass spectra of the tetraquarks $c\bar{c}n\bar{s}$ and $c\bar{c}s\bar{s}$ at varying $R_{0}$ values, as shown in Tables 4-5.
III.2.1 The $c\bar{c}n\bar{n}$ configuration
For the tetraquark state $c\bar{c}n\bar{n}$ with different quantum numbers, the mass spectra and the corresponding meson-meson thresholds are exhibited in Figure 2(a). In this tetraquark system, if the two light quarks are of the same flavor, the isospin of $c\bar{c}n\bar{n}$ is $I=0$ and it is electrically neutral. If one of the two light quarks is a $u$ quark and the other is a $d$ quark, the isospin of $c\bar{c}n\bar{n}$ is $I=1$ and it is a charged particle. At the same time, the possible final state decay channels of these calculated tetraquark states are analyzed according to the quantum numbers and meson thresholds.
There are two $S$-wave hidden-charm tetraquark states with $J^{P(C)}=0^{+(+)}$, whose masses are 3761.1 and 3867.2 MeV respectively. The state with the lowest mass, 3761.1 MeV, lies above the $D\bar{D}$ and $\eta_{c}\pi$ thresholds. Considering the conservation of quantum numbers ($J$, $P$, and $C$), this state can decay naturally into mesons $D\bar{D}$ or $\eta_{c}\pi$. It shows that this state may be observed in the invariant mass spectrum of mesons $D\bar{D}$ or $\eta_{c}\pi$. In addition, in this article, it should be emphasized that we only analyze the $S$-wave decay process of the tetraquark state and ignore the higher-order $P$-wave and $D$-wave decay modes. The reason is that the higher-order decay channels are generally small and are not the main decay modes. The state, 3867.2 MeV, is also above the $\eta_{c}\pi$ and $D\bar{D}$ thresholds but slightly below the $J/\psi\rho$ threshold. Therefore, this state is allowed to decay into mesons $\eta_{c}\pi$, $D\bar{D}$, or $J/\psi\rho$. It can be considered as a candidate for $X(3915)$.
There are three $S$-wave hidden-charm states with $J^{P}=1^{+}$. Their masses are 3754.1, 3872.0, and 3882.3 MeV respectively. this type of tetraquark states can be classified into two categories based on their charge conjugation $C$-parity. For quantum number $J^{PC}=1^{+-}$, the lowest state, 3754.1 MeV, is slightly below the $\eta_{c}\rho$ threshold. It can be considered as a compact tetraquark state. Normally, this state can decay into mesons $\eta_{c}\pi^{+}\pi^{-}$ via an off-shell process for meson $\rho$. In addition, this state can also decay into mesons $J/\psi\pi$. Similarly, the state with a mass of 3882.3 MeV is close to the thresholds $D\bar{D}^{*}$ and $J/\psi\rho$, suggesting that it can easily decay into mesons $D\bar{D}^{*}$ or $J/\psi\rho$. Of course, it can also decay into mesons $\eta_{c}\rho$ or $J/\psi\pi$. For the state with quantum number $J^{PC}=1^{++}$, its mass is very close to the thresholds $J/\psi\rho$ and $D\bar{D}^{*}$. Therefore, this state tends to decay into mesons $J/\psi\rho$ and $D\bar{D}^{*}$. And $J/\psi\rho$ can decay naturally into mesons $J/\psi\pi^{+}\pi^{-}$, which is consistent with previous studies Braaten and Kusunoki (2005). Obviously, it can be concluded that the $|X(3872)\rangle=(c\bar{c}u\bar{u}-c\bar{c}d\bar{d})/\sqrt{2}$ can be considered to be an electrically neutral tetraquark state.
There is an $S$-wave hidden-charm tetraquark state with $J^{P(C)}=2^{+(+)}$. Its mass is 3951.1 MeV, which is above the threshold $J/\psi\rho$. Thus it is possible for the state to decay into mesons $J/\psi\rho$. Again, this state is slightly below the threshold $D^{*}\bar{D}^{*}$, so it may also be found in the $D^{*}\bar{D}^{*}$ invariant mass spectrum.
III.2.2 The $c\bar{c}n\bar{s}$ configuration
The mass spectra of tetraquark $c\bar{c}n\bar{s}$ system are shown in Figure 2(b). For comparison and decay channel analysis, the corresponding meson-meson thresholds are also presented in this figure. For the tetraquark state $c\bar{c}n\bar{s}$, the $C$-parity is undetermined, so the quantum numbers of the $S$-wave states are $J^{P}=0^{+}$, $1^{+}$, and $2^{+}$. In addition, we take $R_{0}=6,8,10$, and $12$ GeV-1, the calculated mass spectra of the $S$-wave tetraquark state $c\bar{c}q\bar{s}$ with varying quantum numbers are listed in Table 4.
There are two $S$-wave tetraquark states with $J^{P}=0^{+}$. Their masses are 3901.1 MeV and 3983.3 MeV, respectively. The lighter state, at 3901.1 MeV, is above the thresholds $\eta_{c}K$ and $D\bar{D}_{s}$, and can easily decay into mesons $\eta_{c}K$ and $D\bar{D}_{s}$. However, due to the large mass gap, the decay widths may be large, making it difficult to observe experimentally in the decay $\eta_{c}K$ and $D\bar{D}_{s}$ channels. The state at 3983.3 MeV can not only decay naturally into the mesons $\eta_{c}K$ and $D\bar{D}_{s}$, but also decay into the mesons $J/\psi K^{*}$. Obviously, this state is close to the threshold $J/\psi K^{*}$, so it is easier to be found in the $J/\psi K^{*}$ invariant mass spectrum.
For $J^{P}=1^{+}$, we find three $S$-wave tetraquark states with masses of 3913.1, 3970, and 3983.2 MeV, respectively. The lightest state, 3913.1 MeV, is above the thresholds $J/\psi K$ and $\eta_{c}K^{*}$, and its decay channels are likely to include $J/\psi K$ or $\eta_{c}K^{*}$. Similarly, the state at 3970 MeV can also decay naturally into the mesons $J/\psi K$ or $\eta_{c}K^{*}$. Of course, this state is very close to the $J/\psi K^{*}$ and $D\bar{D}_{s}^{*}$ thresholds, and it is very likely to decay into the corresponding mesons. The state at 3983.2 MeV is significantly above the thresholds $J/\psi K$and $\eta_{c}K^{*}$, allowing it to decay naturally into the mesons $J/\psi K$and $\eta_{c}K^{*}$. At the same time, this state is located near the thresholds $J/\psi K^{*}$, $D^{*}\bar{D}_{s}$, and $D\bar{D}_{s}^{*}$. There is a high probability that this state will be found in the invariant mass spectrum of $J/\psi K^{*}$, $D^{*}\bar{D}_{s}$, or $D\bar{D}_{s}^{*}$. Obviously, these three hidden-charm tetraquark partner states with strangeness can be candidates for exotic hadrons $Z_{cs}(3985)^{-}$ , $Z_{cs}(4000)^{+}$, or $Z_{cs}(4123)^{-}$.
Table 4: Mass spectra of tetraquark $c\bar{c}n\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)
$M(0^{+})$ $M(1^{+})$ $M(2^{+})$
$R_{0}=6$ 3929, 4011 3942, 3998, 4011 4017
$R_{0}=8$ 3909, 3991 3922, 3978, 3991 3997
$R_{0}=10$ 3901, 3983 3913, 3970, 3983 3989
$R_{0}=12$ 3898, 3980 3910, 3967, 3979 3986
For $J^{P}=2^{+}$, there is an $S$-wave tetraquark state with a mass of 3988.9 MeV. Its mass spectrum lies above the $J/\psi K^{*}$ threshold, allowing it to decay naturally into the mesons $J/\psi K^{*}$. Also, this state is below the threshold $D^{*}\bar{D}_{s}^{*}$, allowing possible decay to the corresponding mesons via off-shell processes. Similarly, this state can also be considered as a candidate for a compact tetraquark state.
III.2.3 The $c\bar{c}s\bar{s}$ configuration
The $S$-wave hidden-charm tetraquark $c\bar{c}s\bar{s}$ system has a well-defined $C$-parity. The possible quantum numbers are $J^{PC}=0^{++}$, $1^{+-}$, and $2^{++}$. Given $R_{0}=10$ GeV-1, the calculated mass spectra of the tetraquark states with the $c\bar{s}c\bar{s}$ configuration are shown in Figure 2(c). At the same time, we also show the corresponding meson-meson thresholds. For varying $R_{0}$, we calculated the mass spectra of the tetraquark $c\bar{c}s\bar{s}$ system under different quantum number assignments, see Table 5.
Table 5: Mass spectra of tetraquark $c\bar{c}s\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)
@@ III.2.3 The $c\bar{c}s\bar{s}$ configuration
$R_{0}=6$ 4060, 4150 4096, 4130 4115 4201
$R_{0}=8$ 4041, 4130 4077, 4111 4096 4182
$R_{0}=10$ 4032, 4122 4069, 4103 4088 4174
$R_{0}=12$ 4030, 4120 4066, 4101 4086 4172
For $S$-wave tetraquark state $c\bar{c}s\bar{s}$ with $J^{PC}=0^{++}$, we find two states with masses of 4032.2 MeV and 4122.3 MeV, respectively. The lowest state, 4032.2 MeV, is above the threshold $D_{s}\bar{D}_{s}$, so its possible decay channel is mesons $D_{s}\bar{D}_{s}$. These results are in good agreement with the experimental observations of the exotic hadron $X(3960)$ by the LHCb collaboration. Therefore, the tetraquark state can be regarded as a good candidate for $X(3960)$. The state at 4122.3 MeV, which lies just below the threshold $D_{s}^{*}\bar{D}_{s}^{*}$, can naturally decay into mesons $D_{s}\bar{D}_{s}$ and $J/\psi\phi$.
There are two $S$-wave tetraquark states with $J^{PC}=1^{+-}$, having masses of 4069.1 and 4102.8 MeV, respectively. The lighter state, 4069.1 MeV, is well above the $\eta_{c}\phi$ threshold, allowing it to naturally decay into mesons $\eta_{c}\phi$. The state at 4102.8 MeV, in addition to the decay mode into mesons $\eta_{c}\phi$, also has a possible decay mode into mesons $D_{s}\bar{D}_{s}^{*}$. It is worth noting that this state is slightly below the threshold $J/\psi\phi$, and a corresponding off-shell decay process may be allowed. Furthermore, there is an $S$-wave tetraquark state with $J^{PC}=1^{++}$ and a mass of 4088.2 MeV. For this state, both $\eta_{c}\phi$ and $D_{s}\bar{D}_{s}^{*}$ decay modes are allowed. This result suggests that the tetraquark state may be a candidate for the exotic hadrons $Y(3940)$ or $X(4140)$.
For $J^{PC}=2^{++}$, we observe an $S$-wave tetraquark state with a mass of 4174.3 MeV. Considering $J^{PC}$ conservation, this tetraquark state can decay into the mesons $J/\psi\phi$. Of course, the channel for the tetraquark state to decay into mesons $D_{s}^{*}\bar{D}_{s}^{*}$ cannot still be prohibited.
III.3 Hidden-bottom tetraquark states with $b\bar{b}n\bar{n}$, $b\bar{b}n\bar{s}$, and $b\bar{b}s\bar{s}$ configurations
Table 6: Mass spectra of tetraquark $b\bar{b}n\bar{n}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)
$M(0^{+(+)})$ $M(1^{+(-)})$ $M(1^{+(+)})$ $M(2^{+(+)})$
$R_{0}=6$ 10583, 10447 10456, 10598 10601 10625
$R_{0}=8$ 10575, 10439 10448, 10591 10593 10618
$R_{0}=10$ 10573, 10437 10446, 10559 10591 10616
$R_{0}=12$ 10572, 10436 10445, 10588 10591 10615
Table 7: Mass spectra of tetraquark $b\bar{b}n\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)
@@
$R_{0}=6$ 10637, 10736 10648, 10727, 10732 10773
$R_{0}=8$ 10622, 10721 10633, 10712, 10717 10758
$R_{0}=10$ 10616, 10715 10626, 10706, 10711 10752
$R_{0}=12$ 10613, 10712 10624, 10704, 10709 10750
Table 8: Mass spectra of tetraquark $b\bar{b}s\bar{s}$ at varying $R_{0}$.
$R_{0}$ (GeV-1) Mass (MeV)